Calculate the mole fraction of each component of a solution containing 65 g of ethanol in 350 g of water.
1
Expert's answer
2020-10-15T03:24:04-0400

"\chi (A)=\frac{n(A)}{n(A) + n(B)} * 100%" %

"\chi (H2O)=\frac{m(H2O)/M(H2O)}{m(H2O)/M(H2O) \ +\ m(C2H5OH)/M(C2H5OH)} * 100%" %

"\chi(H2O) = \frac{350/18}{350/18 \ +\ 65/46} * 100%" % = 93.23%

"\chi(C2H5OH) =100" % - "\chi(H2O)" = 100%-93.23% = 6.77 %